OpenStudy (jazzyfa30):

Solve the equation by using the Square Root Property. x^2 = 10x + 24

OpenStudy (jazzyfa30):

@hartnn

hartnn (hartnn):

ok, so \(x^2-10x=24\) the co-efficient of 'x' here is 10 divide it by 2 and square the result, what u get ?

hartnn (hartnn):

***the co-efficient of 'x' here is -10

OpenStudy (jazzyfa30):

so it would be x^2-5x=24

hartnn (hartnn):

no no.. the co-efficient of x is -10, right ? divide it by 2, you get -5, but then square it, \(\large (-5)^2=...?\)

OpenStudy (jazzyfa30):

2.23

OpenStudy (jazzyfa30):

is that right????

hartnn (hartnn):

umm...noo...how u got thta ?

OpenStudy (jazzyfa30):

i squared 5 isnt that what i was supposed to do

hartnn (hartnn):

5^2 =25

OpenStudy (jazzyfa30):

oh i thought i was supposed to square root it not square my b now what

hartnn (hartnn):

ohh...u took square root...we are supposed to square..

hartnn (hartnn):

ok, we add 25 on both sides \(x^2-10x+25=24+25\) notice that left side is a perfect square! of ?

OpenStudy (jazzyfa30):

49

hartnn (hartnn):

49 is right side, and is the square of 7 what about left side ? x^2-10x+25 is squaree of ?

OpenStudy (jazzyfa30):

idk

hartnn (hartnn):

\((x-5)^2 =... ?\) can u find ?

OpenStudy (jazzyfa30):

x^2- 25

hartnn (hartnn):

nopes, \((a+b)^2=a^2+2ab+b^2\) so, \((x-5)^2 =x^2-10x+25\) so here we have \((x-5)^2=7^2\) take square root on both sides!

OpenStudy (jazzyfa30):

ok

OpenStudy (jazzyfa30):

then what

OpenStudy (jazzyfa30):

im confused

hartnn (hartnn):

did u take square root on both sides ? what u got ?

OpenStudy (jazzyfa30):

idk

hartnn (hartnn):

\((x-5)^2=7^2 \implies x-5 = \pm 7 \\ x-5 =7 \quad \quad x-5=-7\) get 2 values of 'x' from here!

OpenStudy (jazzyfa30):

but 7 isnt apart of any of my answer choices

hartnn (hartnn):

u need to solve x-5 =7 x-5 = -7

OpenStudy (jazzyfa30):

x= 12 x=-2

hartnn (hartnn):

correct!

OpenStudy (jazzyfa30):

ok Solve the equation by completing the square: x2 - 8x + 15 = 0

hartnn (hartnn):

same thing, here, the co-efficient of x is -8 divide it by 2, what u get ?

OpenStudy (jazzyfa30):

-4

hartnn (hartnn):

yes, then square it...what u get ?

OpenStudy (jazzyfa30):

-16

hartnn (hartnn):

-4 *-4 will be +16

OpenStudy (jazzyfa30):

oh

hartnn (hartnn):

so, adding 16 on both sides... \(x^2-8x+16=-15+16\)

OpenStudy (jazzyfa30):

right side:1 left side: idk

OpenStudy (jazzyfa30):

i think its x^2-4x-2x=1

OpenStudy (jazzyfa30):

is that right??????

OpenStudy (jazzyfa30):

@GoldPhenoix please help

hartnn (hartnn):

\(x^2-8x+16 = (x-4)^2\) so, we have \((x-4)^2=1\) just take square root on both sides!

OpenStudy (jazzyfa30):

?????????

hartnn (hartnn):

when you take square root, the square part goes away! (x+4) =+1 or x-4 =-1 find 2 values of x from here...

OpenStudy (jazzyfa30):

(x-3)(x+3) i think

OpenStudy (jazzyfa30):

a. x = + 5 b. x =-3 or x = 5 c. x =3 or x = -5 d. x =3 or x =5

hartnn (hartnn):

x-4 = -1 , x = -1+4 = 3 is correct but for 2nd part, you get x-4 =1 x= 1+4 = 5

hartnn (hartnn):

so, 3,5

OpenStudy (jazzyfa30):

oh yea

OpenStudy (jazzyfa30):

Solve the equation by completing the square: x^2 + 2x = 120

terenzreignz (terenzreignz):

Do you know how to complete the square? :)

OpenStudy (jazzyfa30):

??????

terenzreignz (terenzreignz):

I see. Here's an example:\[\Large x^2 + \color{red}6x =40 \]

terenzreignz (terenzreignz):

What you do is take that coefficient (number beside) the x with NO EXPONENT. In our case, 6.

terenzreignz (terenzreignz):

Take half of it, 3, and then square it, getting 9.

terenzreignz (terenzreignz):

Now, add that to BOTH SIDES of the equation. \[\Large x^2 +\color{red}6x\color{blue}{+9}=40\color{blue}{+9}\]

terenzreignz (terenzreignz):

Simplify. \[\Large x^2+\color{red}6x+9=49\]

terenzreignz (terenzreignz):

You'll notice that the left side is now what you call a 'perfect square trinomial' It may be expressed as... \[\Large (x+\color{red}3)^2=49\]

terenzreignz (terenzreignz):

Notice of course, that 3 is half of the original coefficient of x, which was 6. Now take the square root of both sides: \[\Large \sqrt{(x+3)^2}=\sqrt{49}\] Notice that the square root and the square (exponent 2) cancel out. But the square root of 49 may be either positive or negative 7. \[\Large x+3=\pm7\] Bring the 3 to the other side, you now get \[\Large x = -3\pm7\] which means either x = -3 + 7 = 4 or x = -3 - 7 = -10 Case Closed :)

terenzreignz (terenzreignz):

So... go ahead now, get started with your own problem, using the same concepts. I'll walk you through it if you like :)

OpenStudy (jazzyfa30):

yes please

terenzreignz (terenzreignz):

By your lead :)

OpenStudy (jazzyfa30):

u on skype or call me at 803-937-6271 cus by u typing it it is confusing m please

terenzreignz (terenzreignz):

I don't have skype. come on, we can do this... I'll get you started... \[\Large x^2 +\color{red}2x = 120\]

OpenStudy (jazzyfa30):

then call me i dont understand

OpenStudy (jazzyfa30):

@jessika96